OpenStudy (anonymous):

find the extreme point(s) of the function f(x) = 0.25x^4 + 3x^3 - 18x^2 +10 and classify them

OpenStudy (dumbcow):

what do you mean by extreme points?

OpenStudy (anonymous):

i think they mean extrema, maxima and minima, correct?

OpenStudy (dumbcow):

most likely, in that case differentiate and set equal to zero

OpenStudy (anonymous):

differentiate the equation. then use fishball method

OpenStudy (anonymous):

f(x) = \[0.25x^4 + 3x^3 - 18x^2 +10\] Then differentiate f'(x) = \[x^3 + 9x^2 - 36x\]= \[x(x^2 + 9x - 36)\]= \[x(x+12)(x-3)\]

OpenStudy (anonymous):

Then use fishball method -12 0 3 x - - - 0 + + + x+12 - 0 + + + + + x-3 - - - - - 0 + -------------------------------- - 0 + 0 - 0 + therefore, at x=-12 and x=3 there exists rel. min at x = 0, there exists rel max Hope I helped and I hope I did it right

OpenStudy (dumbcow):

yes i believe this is correct

OpenStudy (anonymous):

Oh gosh thanks so much you guys this really helps